# Echo State Networks for Spatio-Temporal Area-Level Data

### v1

@@ Abstract
Keywords: Areal data; Echo State Network; Graph Convolutional Network; Spatial; Survey.
@@ 1 Introduction
Spatio-temporal area-level datasets are fundamental in official statistics, providing essential data for policy-making and regional planning. For instance, the Bureau of Labor Statistics (BLS) produces the Quarterly Census of Employment and Wages (QCEW) which can be used to analyze business dynamics across the U.S., the Census Bureau produces the Longitudinal Employer-Household Dynamics (LEHD) which can be used to study the flow of jobs and workforces, and the Centers for Disease Control and Prevention (CDC) disseminates numerous spatio-temporal datasets that can be used to track the spread of diseases. Accurately modeling these datasets and generating reliable forecasts can be critical for policymakers to craft informed strategies for future planning. These datasets often exhibit complex temporal dynamics and strong spatial structures, making it crucial to incorporate both temporal and spatial components into the models. Neglecting either aspect can result in inaccurate estimates and potentially misguided policy decisions for data users(Cressie and Wikle,, 2011; Cressie,, 2015; Porter et al.,, 2014, 2015).
@@
A widely used method for handling such datasets is Dynamic Linear Models (DLM), a versatile class of models designed for nonstationary time series that can incorporate various temporal components, such as trend and seasonality (West and Harrison,, 2006; Prado and West,, 2010; Petris et al.,, 2009; Petris,, 2010). However, DLMs are primarily suited for short-lead forecasting. A significant drawback is that they often require substantial computational resources and may encounter numerical issues when applied to large datasets or complex geometries (McDermott and Wikle,, 2017, 2019). Additionally, the modeling process involves making subjective decisions about which state components to include, adding complexity to the model specification.
@@
Recent developments in machine learning and deep learning bring new opportunities in modeling statistical time series. Recurrent Neural Networks (RNN) are highly flexible models capable of capturing complex dynamic patterns through recursive connections between and within neurons (Rumelhart et al.,, 1986; Hochreiter,, 1997). The Echo State Network (ESN) is a specialized type of RNN that efficiently reduces computational requirements by randomly sampling most parameters rather than learning them from data. ESNs have demonstrated success in various statistical time series tasks (Jaeger,, 2002; Lukoševičius,, 2012). For instance, McDermott and Wikle, (2017); Bonas et al., (2024) employed ESNs to forecast Pacific sea surface temperature, McDermott and Wikle, (2019) applied ESNs for soil moisture forecasting, Yoo and Wikle, (2023) utilized ESNs to model wildfire front propagation, and Grieshop and Wikle, (2024) embedded ESNs within a Bayesian hierarchical model to forecast the spread of raccoon rabies. In such applications, spatio-temporal data is often treated as a multivariate time series, with all spatial locations considered as multiple responses, or transformed using empirical orthogonal functions (EOF), with ESN applied to model the EOF coefficients. This approach commonly involves selecting a number of EOFs to retain, which can be determined using the elbow method, setting a threshold for the percentage of variance explained, or by comparing the singular values with those obtained from equivalent uncorrelated data (Hastie et al.,, 2009). However, this general approach may result in the loss of essential information during the dimension reduction process, potentially leading to reduced forecasting performance.
@@
To address this, we aim to capture the neighborhood structure through an embedding similar to Graph Convolutional Networks (GCN). GCNs are specialized convolutional layers in deep learning designed to capture complex features within graph data (Kipf and Welling,, 2016). They are widely regarded as the state-of-the-art method for learning graph representations. Since area-level data can essentially be viewed as an undirected graph, GCN could be well-suited for this task. GCNs operate by averaging the hidden representations of each node with those of its neighbors, and optimizing the target function using gradient-based methods. Although GCNs handle graph features effectively, a key limitation is the significant resource and time requirements for training, as GCNs require the entire graph to be processed in every batch. Various approaches have been proposed to alleviate this issue. For instance, Wu et al., (2019) show that the nonlinear activation in GCNs is not essential, as most of the benefits arise from the local averaging in the GCN layer. They further propose the Simplified Graph Convolutional Network (SGC), which employs a linear activation combined with local averaging among neighbors. Additionally, several authors have explored randomization-based methods to reduce the computational overhead in GCNs. For instance, Huang et al., (2022) demonstrate that employing random weights in GCNs can significantly enhance efficiency while maintaining strong performance in node classification tasks. Liu et al., (2021) combine GCNs with kernel learning using random features, showing promising results in both node classification and clustering tasks.
@@
Building on this idea, we treat the task of handling spatial features as a problem of learning good representations from data (Bengio et al.,, 2013). Specifically, we draw inspiration from the expanded representation concept proposed by Sutton et al., (2014), which involves a two-layer model: the first layer expands the original input into a high-dimensional feature space, while the second layer maps this representation to the output. Classic models such as Rosenblatt’s perceptron and modern deep learning frameworks can all be seen as instances of expanded representation learning. Notably, Gallant, (1987) and Sutton et al., (2014) advocate the use of a random fixed layer for input expansion, a technique known as random representation. This method efficiently handles high-dimensional inputs and enables faster computation, making it well-suited for generating input embeddings for ESN.
@@
In this paper, we propose a novel approach for capturing spatio-temporal dynamics often found in areal data and apply the approach to data disseminated by Eurostat. We employ an ensemble ESN with a specialized input embedding designed for areal data, integrating concepts from Random Representation and SGC layer. Our spatial input embedding leverages the local averaging properties of the SGC, and uses the fixed randomly sampled parameters to effectively capture spatial features. By addressing spatial dependencies at the input stage, this approach significantly enhances ESN performance while maintaining its computational efficiency. The paper is organized as follows. Section 2 reviews the concepts of ESN and GCN. In Section 3, we describe our areal input embedding, along with its uncertainty quantification and hyperparameter selection. Section 4 provides an illustration of forecasting performance on the Eurostat’s night spent at tourism accommodations dataset. Finally, we provide a brief summary in Section 5.
@@ 2.1 Echo State Network
As described in McDermott and Wikle, (2017), the fundamental architecture of an ESN is described as follows
[…]
@@
[…]
$\bm{W}_{in}=[w_{i,j}]_{i,j}:w_{i,j}=\gamma_{i,j}^{in}\text{Unif}(-a_{in},a_{in})+(1-\gamma_{i,j}^{in})\delta_{0},$
@@
where $\bm{Y}_{t}\in\mathbb{R}^{n_{y}}$, $\bm{h}_{t}\in\mathbb{R}^{n_{h}}$, and $\bm{x}_{t}\in\mathbb{R}^{n_{x}}$ represent the output vector, the hidden state vector, and the input vector at time $t$, respectively. The matrices $\bm{W}_{out}\in\mathbb{R}^{n_{y}\times n_{h}}$, $\bm{W}_{res}\in\mathbb{R}^{n_{h}\times n_{h}}$, and $\bm{W}_{in}\in\mathbb{R}^{n_{h}\times n_{x}}$ denote the output weights, the hidden weights, and the input weights at time $t$, respectively. $\delta_{0}$ denotes the Dirac function centered at zero, $\alpha$ is the leaking rate from the previous reservoir state, $\nu$ is the scaling factor for $\bm{W}_{res}$, and $\lambda_{w}$ is the largest eigenvalue of $\bm{W}_{res}$.
The hidden layer, often known as the reservoir computing layer, plays a crucial role in modeling temporal dynamics. The reservoir itself is constructed using the randomly sampled weight matrix, $\bm{W}_{res}$, which is scaled by its spectral radius, $|\lambda_{w}|$. This process creates a large pool of nonlinear transformations of the input signals, combining the scaled input data from the current time with outputs from the previous reservoir states. These transformations, or hidden states, are designed to capture the underlying dynamics of the time series.
Once the hidden states are generated, the output weights $\bm{W}_{out}$, also called readout weights, are trained to map these states to the desired output. Notably, the only parameter that requires training in this approach is the readout weight, which can be efficiently estimated using simple ridge regression. The loss function for this estimation is given by $L=\frac{1}{N}\sum_{t=1}^{T}\left((\bm{Y}_{t}-\bm{W}_{out}\bm{h}_{t})(\bm{Y}_{t}-\bm{W}_{out}\bm{h}_{t})^{\prime}\right)+\tau\sum_{j=1}^{n_{y}}\|\bm{W}_{out}^{2}\|_{2}$, where $\tau$ is the penalty term. As a result, reservoir computing only requires a minimal amount of computation resources.
To ensure optimal performance, the hyperparameters of the ESN must be fine-tuned. Specifically, we need to determine the best values for the reservoir size, $\bm{W}_{res}$, denoted as $n_{h}$, the scaling factor of $\bm{W}_{res}$, $\nu$, the distribution parameters of the reservoir and input weights, the leaking rate $\alpha$ from the previous reservoir state, and the penalty coefficient in ridge regression, $\tau$. In practice, particularly in spatio-temporal modeling applications, the most influential hyperparameters are $\nu$, $n_{h}$, and $\tau$ (McDermott and Wikle,, 2017; Yoo and Wikle,, 2023).
@@ 2.2 Graph Convolutional Network
Kipf and Welling, (2016) introduced the GCN to learn feature representations for each node in a graph. Given a graph $\mathcal{G}$ with $n$ nodes, each having $p$ features, and a proximity matrix $\bm{A}\in\mathbb{R}^{n\times n}$, let $\bm{H}_{0}=\bm{X}\in\mathbb{R}^{n\times p}$ represent the input feature matrix. For $i>1$, let $\bm{H}_{i}\in\mathbb{R}^{n\times h_{i}}$ denote the feature matrix learned at the $i$th graph convolutional layer. $\bm{H}_{i}$ is updated as
@@
where $\bm{\Theta}_{i-1}\in\mathbb{R}^{h_{i-1}\times h_{i}}$ is the weight matrix at the (i-1)th layer, $\widetilde{\bm{A}}=\bm{A}+\bm{I}\in\mathbb{R}^{n\times n}$, which denotes the proximity matrix with self-loops included, and $\widetilde{\bm{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix with entries given by the row sums of $\widetilde{\bm{A}}$. This definition of the proximity matrix is beneficial because it allows for handling isolated nodes, where the node’s features consist solely of its own attributes rather than incorporating features from neighboring nodes.
Kipf and Welling, (2016) further show that the update rule of GCNs can be viewed as a first-order approximation of localized spectral filters on graphs. Let $\bm{L}\in\mathbb{R}^{n\times n}$ denotes the normalized graph Laplacian of $\mathcal{G}$, $\bm{U}\in\mathbb{R}^{n\times n}$ be the eigenvectors of $\bm{L}$, and $\bm{\Lambda}\in\mathbb{R}^{n\times n}$ be the diagonal matrix of eigenvalues of $\bm{L}$. We can write $\bm{L}$ as
@@
where $\bm{A}$ is defined as in (1) and $\bm{D}$ is the corresponding degree matrix. The spectral convolution operation on graph can be defined as
@@
where $\bm{x}$ is the input signal on the nodes, and $\bm{g}_{\theta}$ is a spectral filter parameterized by $\theta$ in the Fourier domain. Kipf and Welling, (2016) approximate $\bm{g}_{\theta}$ using first-order Chebyshev polynomials $\bm{g}_{\theta}\approx\theta_{0}+\theta_{1}\bm{\Lambda}$ (Hammond et al.,, 2011). Thus, (3) can be effectively rewritten as
@@
To address overfitting and reduce the number of parameters, Hammond et al., (2011) further set $\theta=2\theta_{0}=-\theta_{1}$, and allow the parameter $\theta$ to be shared across the entire graph. Additionally, to avoid numerical instabilities, they redefine the adjacency and degree matrices as $\widetilde{\bm{A}}=\bm{A}+\bm{I}$ and $\widetilde{\bm{D}}_{ii}=\sum_{j}\widetilde{\bm{A}}_{ij}$, respectively. We thus obtain the final form of the GCN updating rule as shown in (1).
Kipf and Welling, (2016) consider applying a nonlinear activation function to the output of (1) to make GCNs more flexible. However, Wu et al., (2019) show that the ability for feature extraction on graphs is due to the local averaging introduced by (1), and that the nonlinear activation is not necessary. They denote the GCN with a linear activation function as the Simple Graph Convolution (SGC) and demonstrate that it corresponds to a fixed low-pass filter on the graph spectral domain.

### v2

@@ Abstract
Keywords: Areal data; Bayesian; Echo State Network; Graph Convolutional Network; Spatial; Survey.
@@ 1 Introduction
Spatio-temporal area-level datasets are fundamental in official statistics, providing essential data for policy-making and regional planning. For instance, the Bureau of Labor Statistics (BLS) produces the Quarterly Census of Employment and Wages (QCEW) which can be used to analyze business dynamics across the U.S., the Census Bureau produces the Longitudinal Employer-Household Dynamics (LEHD) which can be used to study the flow of jobs and workforces, and the Centers for Disease Control and Prevention (CDC) disseminates numerous spatio-temporal datasets that can be used to track the spread of diseases. Accurately modeling these datasets and generating reliable forecasts can be critical for policymakers to craft informed strategies for future planning. These datasets often exhibit complex temporal dynamics and strong spatial structures, making it crucial to incorporate both temporal and spatial components into the models. Neglecting either aspect can result in inaccurate estimates and potentially misguided policy decisions for data users(cressie2011statistics; cressie2015statistics; porter2014spatial; porter2015small).
@@
A widely used method for handling such datasets is Dynamic Linear Models (DLM), a versatile class of models designed for nonstationary time series that can incorporate various temporal components, such as trend and seasonality (west2006bayesian; prado2010time; dlm_book; dlm_r). However, DLMs are primarily suited for short-lead forecasting. A significant drawback is that they often require substantial computational resources and may encounter numerical issues when applied to large datasets or complex geometries (mcdermott2017ensemble; mcdermott2019deep). Additionally, the modeling process involves making subjective decisions about which state components to include, adding complexity to the model specification.
@@
Recent developments in machine learning and deep learning bring new opportunities in modeling statistical time series. For example, eliwa2024optimal uses the adaptive network-based fuzzy inference system to predict gasoline price. hassan2024optimizing uses language models to automatically predict diseases from their symptoms. Among all machine learning models, Recurrent Neural Networks (RNN) are highly flexible models capable of capturing complex dynamic patterns through recursive connections between and within neurons (rumelhart1986learning; hochreiter1997long). The Echo State Network (ESN) is a specialized type of RNN that efficiently reduces computational requirements by randomly sampling most parameters rather than learning them from data. ESNs have demonstrated success in various statistical time series tasks (jaeger2002adaptive; lukovsevivcius2012practical). For instance, mcdermott2017ensemble; bonas2024calibrated employed ESNs to forecast Pacific sea surface temperature, mcdermott2019deep applied ESNs for soil moisture forecasting, yoo2023using utilized ESNs to model wildfire front propagation, and grieshop2024echo embedded ESNs within a Bayesian hierarchical model to forecast the spread of raccoon rabies. wang2025hierarchical extended ESNs to accommodate count data and applied them to forecast the number of graduate students across universities in the United States. In such applications, spatio-temporal data is often treated as a multivariate time series, with all spatial locations considered as multiple responses, or transformed using empirical orthogonal functions (EOF), with ESN applied to model the EOF coefficients. This approach commonly involves selecting a number of EOFs to retain, which can be determined using the elbow method, setting a threshold for the percentage of variance explained, or by comparing the singular values with those obtained from equivalent uncorrelated data (hastie2009elements). However, this general approach may result in the loss of essential information during the dimension reduction process, potentially leading to reduced forecasting performance.
@@
To address this, we aim to capture the neighborhood structure through an embedding similar to Graph Convolutional Networks (GCN). GCNs are specialized convolutional layers in deep learning designed to capture complex features within graph data (kipf2016semi). They are widely regarded as the state-of-the-art method for learning graph representations. Since area-level data can essentially be viewed as an undirected graph, GCN could be well-suited for this task. GCNs operate by averaging the hidden representations of each node with those of its neighbors, and optimizing the target function using gradient-based methods. Although GCNs handle graph features effectively, a key limitation is the significant resource and time requirements for training, as GCNs require the entire graph to be processed in every batch. Various approaches have been proposed to alleviate this issue. For instance, wu2019simplifying show that the nonlinear activation in GCNs is not essential, as most of the benefits arise from the local averaging in the GCN layer. They further propose the Simplified Graph Convolutional Network (SGC), which employs a linear activation combined with local averaging among neighbors. Additionally, several authors have explored randomization-based methods to reduce the computational overhead in GCNs. For instance, huang2022graph demonstrate that employing random weights in GCNs can significantly enhance efficiency while maintaining strong performance in node classification tasks. liu2021random combine GCNs with kernel learning using random features, showing promising results in both node classification and clustering tasks.
@@
Building on this idea, we treat the task of handling spatial features as a problem of learning good representations from data (bengio2013representation). Specifically, we draw inspiration from the expanded representation concept proposed by sutton2014online, which involves a two-layer model: the first layer expands the original input into a high-dimensional feature space, while the second layer maps this representation to the output. Classic models such as Rosenblatt’s perceptron and modern deep learning frameworks can all be seen as instances of expanded representation learning. Notably, gallant1987random and sutton2014online advocate the use of a random fixed layer for input expansion, a technique known as random representation. This method efficiently handles high-dimensional inputs and enables faster computation, making it well-suited for generating input embeddings for ESN.
@@
In this paper, we propose a novel approach for capturing spatio-temporal dynamics often found in areal data and apply the approach to data disseminated by Eurostat. We employ an ensemble ESN with a specialized input embedding designed for areal data, integrating concepts from Random Representation and SGC layer. We refer to the ESN with areal random representation at the input stage as AESN. Our spatial input embedding leverages the local averaging properties of the SGC, and uses the fixed randomly sampled parameters to effectively capture spatial features. By addressing spatial dependencies at the input stage, this approach significantly enhances ESN performance while maintaining its computational efficiency.
We demonstrate the performance of the proposed method using Eurostat’s tourism occupancy data. Tourism plays a crucial role in shaping regional economies, and tourism occupancy is a valuable indicator for the popularity of the tourism industry in Europe. It is essential to understand the spatio-temporal patterns of tourism occupancy and accurately forecast them into the future. However, modeling this data poses significant challenges due to its inherent complexity. As an area-level survey dataset, tourism occupancy data exhibits intricate spatial patterns within the European NUTS  [^5] Nomenclature of Territorial Units for Statistics (NUTS) is the official classification of subregions used by Eurostat. level 2 regions, coupled with complex temporal dynamics. These characteristics make the proposed method suitable for capturing the underlying patterns.
The paper is organized as follows. Section 2 reviews the concepts of ESN and GCN. In Section 3, we describe our areal input embedding, along with its uncertainty quantification and hyperparameter selection. Section 4 provides an illustration of forecasting performance on the Eurostat’s night spent at tourism accommodations dataset. Finally, we provide a brief summary in Section 5.
@@ 2.1 Echo State Network
As described in mcdermott2017ensemble, the fundamental architecture of an ESN is described as follows
$\text{output layer}:\qquad$ $\bm{Y}_{t}=\bm{W}_{out}\bm{h}_{t}+\bm{\epsilon}_{t},\bm{\epsilon}_{t}\sim N(\bm{0},\sigma^{2}\bm{I}),$
$\text{hidden layer}:\qquad$ $\bm{h}_{t}=(1-\alpha)\bm{h}_{t-1}+\alpha\widetilde{\bm{h}}_{t},$
@@
$\text{parameters}:\qquad$ $\bm{W}_{res}=[w_{i,l}]_{i,l}:w_{i,l}=\gamma_{i,l}^{res}\text{Unif}(-a_{res},a_{res})+(1-\gamma_{i,l}^{res})\delta_{0},$ (1)
$\bm{W}_{in}=[w_{i,j}]_{i,j}:w_{i,j}=\gamma_{i,j}^{in}\text{Unif}(-a_{in},a_{in})+(1-\gamma_{i,j}^{in})\delta_{0},$ (2)
@@
where $\bm{Y}_{t}\in\mathbb{R}^{n_{y}}$, $\bm{h}_{t}\in\mathbb{R}^{n_{h}}$, and $\tilde{\bm{x}_{t}}\in\mathbb{R}^{n_{x}}$ represent the output vector, the hidden state vector, and the input vector at time $t$, respectively. The matrices $\bm{W}_{out}\in\mathbb{R}^{n_{y}\times n_{h}}$, $\bm{W}_{res}\in\mathbb{R}^{n_{h}\times n_{h}}$, and $\bm{W}_{in}\in\mathbb{R}^{n_{h}\times n_{x}}$ denote the output weights, the hidden weights, and the input weights at time $t$, respectively. $\delta_{0}$ denotes the Dirac function centered at zero, $\alpha$ is the leaking rate from the previous reservoir state, $\nu$ is the scaling factor for $\bm{W}_{res}$, and $\lambda_{w}$ is the largest eigenvalue of $\bm{W}_{res}$.
The hidden layer, often known as the reservoir computing layer, plays a crucial role in modeling temporal dynamics. The reservoir itself is constructed using the randomly sampled weight matrix, $\bm{W}_{res}$, which is scaled by its spectral radius, $|\lambda_{w}|$. This process creates a large pool of nonlinear transformations of the input signals, combining the scaled input data from the current time with outputs from the previous reservoir states. These transformations, or hidden states, are designed to capture the underlying dynamics of the time series. A fundamental aspect of ESNs is their reliance on the echo state property (ESP), which states that the reservoir state is driven by the past input sequence, rather than its initial state (jaeger2002adaptive; lukovsevivcius2012practical). In practice, the ESP is typically maintained by constraining the spectral radius of the reservoir weights to be less than one (jaeger2002adaptive; lukovsevivcius2012practical).
Once the hidden states are generated, the output weights $\bm{W}_{out}$, also called readout weights, are trained to map these states to the desired output. Notably, the only parameter that requires training in this approach is the readout weights, which can be efficiently estimated using simple ridge regression. The loss function for this estimation is given by $L=\frac{1}{N}\sum_{t=1}^{T}\left((\bm{Y}_{t}-\bm{W}_{out}\bm{h}_{t})(\bm{Y}_{t}-\bm{W}_{out}\bm{h}_{t})^{\prime}\right)+\tau\sum_{j=1}^{n_{y}}\|\bm{W}_{out}^{2}\|_{2}$, where $\tau$ is the penalty term. As a result, reservoir computing only requires a minimal amount of computation resources.
To ensure optimal performance, the hyperparameters of the ESN must be fine-tuned. Specifically, we need to determine the best values for the reservoir size, $\bm{W}_{res}$, denoted as $n_{h}$, the scaling factor of $\bm{W}_{res}$, $\nu$, the distribution parameters of the reservoir and input weights, the leaking rate $\alpha$ from the previous reservoir state, and the penalty coefficient in ridge regression, $\tau$. In practice, particularly in spatio-temporal modeling applications, the most influential hyperparameters are $\nu$, $n_{h}$, and $\tau$ (mcdermott2017ensemble; yoo2023using).
@@ 2.2 Graph Convolutional Network
kipf2016semi introduced the GCN to learn feature representations for each node in a graph. Given a graph $\mathcal{G}$ with $n$ nodes, each having $p$ features, and a proximity matrix $\bm{A}\in\mathbb{R}^{n\times n}$, let $\bm{H}_{0}=\bm{X}\in\mathbb{R}^{n\times p}$ represent the input feature matrix. For $i>1$, let $\bm{H}_{i}\in\mathbb{R}^{n\times h_{i}}$ denote the feature matrix learned at the $i$th graph convolutional layer. $\bm{H}_{i}$ is updated as
@@
where $\bm{\Theta}_{i-1}\in\mathbb{R}^{h_{i-1}\times h_{i}}$ is the weight matrix at the (i-1)th layer, $\widetilde{\bm{A}}=\bm{A}+\bm{I}\in\mathbb{R}^{n\times n}$, which denotes the proximity matrix with self-loops, and $\widetilde{\bm{D}}\in\mathbb{R}^{n\times n}$ is the diagonal degree matrix with entries given by the row sums of $\widetilde{\bm{A}}$. This definition of the proximity matrix is beneficial because it allows for handling isolated nodes, where the node’s features consist solely of its own attributes rather than incorporating features from neighboring nodes.
kipf2016semi further show that the update rule of GCNs can be viewed as a first-order approximation of localized spectral filters on graphs. Let $\bm{L}\in\mathbb{R}^{n\times n}$ denotes the normalized graph Laplacian of $\mathcal{G}$, $\bm{U}\in\mathbb{R}^{n\times n}$ be the eigenvectors of $\bm{L}$, and $\bm{\Lambda}\in\mathbb{R}^{n\times n}$ be the diagonal matrix of eigenvalues of $\bm{L}$. We can write $\bm{L}$ as
@@
where $\bm{A}$ is defined as in GCN and $\bm{D}$ is the corresponding degree matrix. The spectral convolution operation on graph can be defined as
@@
where $\bm{x}$ is the input signal on the nodes, and $\bm{g}_{\theta}$ is a spectral filter parameterized by $\theta$ in the Fourier domain. kipf2016semi approximate $\bm{g}_{\theta}$ using first-order Chebyshev polynomials $\bm{g}_{\theta}\approx\theta_{0}+\theta_{1}\bm{\Lambda}$ (hammond2011wavelets). Thus, (5) can be effectively rewritten as
@@
To address overfitting and reduce the number of parameters, hammond2011wavelets further set $\theta=2\theta_{0}=-\theta_{1}$, and allow the parameter $\theta$ to be shared across the entire graph. Additionally, to avoid numerical instabilities, they redefine the adjacency and degree matrices as $\widetilde{\bm{A}}=\bm{A}+\bm{I}$ and $\widetilde{\bm{D}}_{ii}=\sum_{j}\widetilde{\bm{A}}_{ij}$, respectively. We thus obtain the final form of the GCN updating rule as shown in (3).
kipf2016semi consider applying a nonlinear activation function to the output of (3) to make GCNs more flexible. However, wu2019simplifying show that the ability for feature extraction on graphs is due to the local averaging introduced by (3), and that the nonlinear activation is not necessary. They denote the GCN with a linear activation function as the Simple Graph Convolution (SGC) and demonstrate that it corresponds to a fixed low-pass filter on the graph spectral domain.